### v2

# From Multiple-Choice to Extractive QA: A Case Study for English and Arabic

Abstract
The rapid evolution of Natural Language Processing (NLP) has favoured major languages such as English, leaving a significant gap for many others due to limited resources. This is especially evident in the context of data annotation, a task whose importance cannot be underestimated, but which is time-consuming and costly. Thus, any dataset for resource-poor languages is precious, in particular when it is task-specific. Here, we explore the feasibility of repurposing an existing multilingual dataset for a new NLP task: we repurpose a subset of the Belebele dataset  Bandarkar et al. (2023), which was designed for multiple-choice question answering (MCQA), to enable the more practical task of extractive QA (EQA) in the style of machine reading comprehension. We present annotation guidelines and a parallel EQA dataset for English and Modern Standard Arabic (MSA). We also present QA evaluation results for several monolingual and cross-lingual QA pairs including English, MSA, and five Arabic dialects. We aim to help others adapt our approach for the remaining 120 Belebele language variants, many of which are deemed under-resourced. We also provide a thorough analysis and share insights to deepen understanding of the challenges and opportunities in NLP task reformulation.
From Multiple-Choice to Extractive QA:
A Case Study for English and Arabic
Teresa Lynn,1 Malik H. Altakrori,2 Samar M. Magdy1 Rocktim Jyoti Das,1 Chenyang Lyu,1 Mohamed Nasr,2 Younes Samih,2 Kirill Chirkunov,1 Alham Fikri Aji,1 Preslav Nakov,1 Shantanu Godbole,2 Salim Roukos,2 Radu Florian,2 Nizar Habash1,3 1Mohamed bin Zayed University of Artificial Intelligence 2IBM Research AI  3New York University Abu Dhabi teresa.lynn@mbzuai.ac.ae, malik.altakrori@ibm.com, nizar.habash@nyu.edu
1 Introduction
Recent years have brought about very fast developments in Natural Language Processing (NLP). However, this progress has not been equal for all languages, and most research has focused on English and a handful of other languages, with the vast majority of other languages being overlooked Joshi et al. (2020); Rehm and Way (2023). This is mainly due to the lack of data resources, which hampers the development of NLP tools for these languages. While many resources have been developed for some languages, they are often for very specific tasks, and for very specific formulations of these tasks.
Starting from a motivation of creating cross-lingual Question Answering (QA) datasets for an under-resourced pair, Dialectal Arabic (DA) – Modern Standard Arabic (MSA), we explore the possibility of re-purposing the multilingual Belebele dataset Bandarkar et al. (2023), which was created for the task of multiple-choice question answering, to be suitable for extractive question answering in the style of machine reading comprehension.
In the field of Question Answering, there are a number of different ways to answer a question  Wang (2022), e.g., (a) multiple-choice QA (MCQA) with answer choices, and context provided, (b) extractive QA (EQA) with context provided, (c) abstractive QA with context provided, and (d) open-ended QA with no context provided. An interesting research question is whether we can traverse between them, i.e., if we have a resource for one QA task, can we repurpose it easily for a different reformulation of the task? Here we study this research question, in the context of transferring from MCQA to EQA, as we deem EQA more useful in real-world applications where the correct answer is often not known. Our contributions are:
  • We explore the possibility for repurposing a MCQA dataset as an EQA dataset.
  • We create and release a new parallel EQA dataset for MSA and English (EN).
  • We provide a dialectal Arabic QA benchmark by evaluating our EQA system’s performance on monolingual and cross-lingual QA pairs of {EN, MSA}–{EN, MSA, five DAs}.
  • We provide guidelines and a strong foundation for creating EQA datasets for the remaining 120 Belebele languages.
  • We perform careful analysis and discuss the lessons learned as a guide for future research in repurposing NLP datasets.
2 Related Work
2.1 Multilingual QA Datasets
Numerous multilingual QA datasets have already been developed e.g., XQUAD Artetxe et al. (2020), TyDi QA Clark et al. (2020), and MLQA Lewis et al. (2020), which focus on extractive-based QA. The recent ArabicaQA dataset Abdallah et al. (2024) represents a significant leap forward, being the first large-scale dataset specifically designed for Arabic QA: it has over 89k questions, derived from a diverse set of Standard Arabic documents, providing a robust platform for advancing Arabic NLP, particularly in the context of large language models. Additionally, alternative formats of multilingual QA have been explored, including open-ended QA with Mintaka Sen et al. (2022) and MCQA with Belebele Bandarkar et al. (2023). With the exception of Belebele, these datasets encompass a relatively narrow range of languages and do not include Arabic dialects. To address this gap, we propose leveraging the Belebele dataset, to enable the inclusion of its dialectal content into a new parallel DA – MSA/EN dataset.
2.2 Converting MCQA Datasets
Some attempts have been made to convert existing datasets into a different format. Specifically for QA, there has been conversion of MCQA datasets into natural language inference (NLI) Demszky et al. (2018); Khot et al. (2018): the premise is taken from the context paragraph, the correct choice is used as an entailment label, and the other choices are used as neutral or contradictory. Both approaches require effort to rephrase the question into a statement, e.g., using rule-based methods Demszky et al. (2018) or human annotation Khot et al. (2018). Hadifar et al. (2023) introduced a new dataset specifically designed for the educational setting, focusing on converting between multiple-choice and open-ended question formats. Automatically generating MCQA has also been explored Mitkov and Ha (2003); Kurdi et al. (2020); Ai et al. (2015); Karamanis et al. (2006), where the challenge is to generate distractor choices.
To our knowledge, there is no previous work on converting MCQA to EQA datasets. In this paper, we investigate the feasibility of such a task with an eye towards future automation of the process.
Table 1: English examples demonstrating some interesting types of MCQs in the Belebele dataset, and the respective issues they pose for extractive QA in this work. The Arabic version is shown in Appendix C.
2.3 A Note on Arabic and its Dialects
The Arabic language is a family of variants, among which Modern Standard Arabic (MSA) is the Arab world’s shared language of culture, media, and education. However, MSA is not the native language of any speaker of Arabic, whose day-to-day language is typically a local variety, i.e., Dialectal Arabic (DA) that can be quite different from MSA and other dialects Salameh et al. (2018). MSA and DA coexist in what is called a diglossic relationship where different variants are used in different contexts Ferguson (1959). In the context of Arabic QA, there are many efforts that focus primarily on MSA Shaheen and Ezzeldin (2014); Mozannar et al. (2019); Biltawi et al. (2021); Alwaneen et al. (2022). A recent exception is the work of Faisal et al. (2021), which works on spoken dialectal QA and includes Arabic among other languages.
The common wisdom in Arabic NLP is that systems should be robust to handling DA forms, but should generate output primarily in MSA. This is reasonable for QA in particular, given the relatively larger trusted content in MSA compared to DA which dominates social media. As such, in this paper, we focus on modeling QA that accepts questions in a range of dialects and provides answers in standard languages (MSA and EN). We selected the following five representative dialects from across the Arab world to work with: Egyptian (arz), Iraqi (acm), Moroccan (ary), Gulf (Najdi; ars), and Levantine (North; apc).
3 MCQA-to-EQA Conversion Challenges
3.1 Belebele Dataset
Belebele Bandarkar et al. (2023) is an open-source multiple-choice machine reading comprehension (MRC) dataset that covers 122 languages and language variants, including English, MSA, and five Arabic dialects. Each question is based on a short passage from the Flores-200 dataset Team et al. (2022) and has four multiple-choice answers, with the correct answer identified. This benchmark dataset enabled the evaluation of natural language understanding (NLU) across high/medium/low-resource languages and language variants.
With respect to the style of the dataset content, the creators of the Belebele dataset indicated that they selected multiple-choice questions (MCQs) because it would lead to the fairest evaluation across languages and MCQs enable the ability to scale to many languages when translating from English. They worked with a Language Service Provider to create the questions and the multi-choice answers in an iterative process, by developing guidelines that included rules such as no use of double negatives, only use answers that are decently plausible to require the test-taker to fully read and understand the passage Bandarkar et al. (2023).
Our effort is part of a larger project on Arabic QA; as such we deemed the Belebele dataset to be a suitable starting point for establishing an EQA benchmark for Arabic dialect QA when paired with MSA and English passages. Our five dialects of focus represent the full spectrum of variations across the Arab world: Gulf, Levantine, Iraqi, Egyptian, and Moroccan. We have questions in English, MSA and Arabic dialects, with the answers in MSA and English only.
Broadly defined, our first task in the conversion from MCQA to EQA is to identify the spans in the passages that answer the MCQs.
3.2 The MCQ Zoo
When examining the different Belebele MCQs and their associated passages, we recognized a considerable number of MCQ types that vary on a continuum from simple word matching to required passage and world understanding. Table 1 presents a set of English examples along this continuum to highlight the challenges that hindered us from employing a fully automated process in identifying text spans in the associated passages. The Arabic version is in Appendix C.
The simplest MCQ types are perhaps Exact Match, Partial Overlap Match and Morphological Match, Table 1.(a, b, c), respectively, where the MCQ answer is “present” in the passage. Relatively more complex are cases that may be resolvable through Paraphrasing, Entailment, or Semantic Similarity, in Table 1.(d, e, f), respectively. Next comes the cases that require higher levels of awareness of context, from that of Coreference in the text all the way to Pragmatic Knowledge, in Table 1.(g-i). Finally we present examples of complex kinds of MCQ, where wording of the question is highly dependent on the provided multi-choice answers. While a span can be identified in principle in Table 1.(j), the question may not stand alone with the passage in Table 1.(k, l).
These challenges made it apparent that (a) not all MCQAs are usable for the EQA task; and (b) that human annotation was a more realistic approach than automating the conversion of the dataset. While the first few MCQ types suggest that an automatic process could be used to identify the QA span, the rest, in increasingly difficulty, show that the MCQ answers are not as helpful. The last two types are not usable in our assessment for this task. We took all of these insights into consideration as we developed the guidelines and annotated the spans, which we discuss next.
4 Annotation Process
In this study, annotation refers to marking the appropriate ‘answer span’ in the original passage in response to a question, or labelling a question as unanswerable (‘X’), as shown in Table 1. Our annotations focused on EN and MSA parallel passages from the Belebele dataset and we call our annotated set Belebele-EQA. Our repurposing approach involved the following steps.
4.1 Automatic Filtering
The original Belebele dataset contains 900 question-answer pairs. However, as highlighted in Section 3.2, many of these multiple-choice questions did not suit the EQA task. As a first pass, we automatically removed QA pairs that were deemed unsuitable based on the following set of keywords: “Which of the following”, “Select all that apply", “Choose the correct”, “According to the passage”, “Based on the information given”. Given that the dataset is parallel, we carried out this automatic filtering on the English content first and subsequently removed the same QA pairs in the corresponding MSA content. After this first filtering pass, we were left with 415 QA pairs, whose EQA answer spans we annotated manually.
4.2 Pilot study
We first carried out a pilot annotation study in order to fully understand the nature of the Belebele dataset and develop our annotation guidelines. The pilot annotation study was carried out on the first 100 QAs for both EN and MSA. The annotators were presented with the Belebele passage, the original Belebele question, the ‘correct answer’ from the multiple-choice answer list [^1] Available from the Belebele data release https://github.com/facebookresearch/belebele and a column in which to write the span. Two native Arabic speakers annotated 50 MSA QAs each, and two fluent English speakers each annotated 50 EN QAs. It was during this pilot study that the extent of the challenges highlighted in Section 3 became known.
4.3 Annotation Guidelines
Informed by the findings of the initial pilot study, we created two separate versions of the guidelines for English and Arabic, accounting for linguistic differences that influenced the wording of the instructions and the need for language-specific examples. See Appendices D and E for the complete EN and MSA guidelines, respectively.
Also informed by the pilot study, the original multiple choice answer was hidden during the annotation, as it had been deemed to be more distracting than helpful in the pilot round (see discrepancies between MC answers and spans in Table 1).
As part of the annotation task, the annotators would mark the span in a copy of the passage using \$!(text)!\$ as a span delimiter, while keeping the span as short as possible – in the style of SQUAD Rajpurkar et al. (2016).
Annotators were also given clear guidance on what to do with articles, prepositions, punctuation, and instances where the answer appeared more than once in the passage (Appendices D and E). Equipped with better awareness of the issues described in Section 3.2, for each QA pair, the annotators could either: (i) annotate the span and if necessary, label the QA as ‘BB’ to indicate it as being complex and problematic, as per Table 1.(d,e,g,i) or (ii) label the QA as ‘X’ to be removed, due to the answer being impossible to find, as per Table 1.(k,l). Finally, annotators were advised that it was more important to find a relevant span (in which evidence of the answer could be found) than finding an exact span that may not exist. This guidance was a significant help given the many differences between the nature of MCQA and EQA.
4.4 English Data Annotation
We took a bootstrapping approach to annotation across the two languages. Given the parallel nature of the dataset, the EN annotation was carried out first, followed by a semi-automatic approach to annotating the MSA data (Section 4.5). The annotations were done on the remaining 415 QA pairs in the Belebele dataset (including a review and correction of the first 100 from the pilot round). Eleven annotators, both fluent and native English speakers, annotated the QA pairs, with the annotations subsequently reviewed by independent reviewers.
The distribution of the QA pairs was as follows: 50 QA pairs were used in our IAA study (see Section 4.6) and were annotated by three annotators. The remaining 365 pairs were distributed amongst the other eight annotators, where six annotators had 50 pairs each, and the remaining two annotators had 42 and 23, respectively. The annotators were given the English Belebele passage, the original question and a copy of the passage to mark the answer span. They could also add comments for problematic instances that required broader discussion or verification.
As a result of the annotation process, 86 out of the 415 QA pairs were labelled as ‘X’ and excluded. Out of the remaining 329 pairs, 44 were labelled as ‘BB’. These BB QAs remain part of the final dataset, as we are also interested in establishing what difficulties they pose for a QA system, given the need for reasoning and resolving linguistic phenomena like paraphrasing, pragmatics and coreference resolution.
4.5 Arabic Data Annotation
For the MSA data, we exploited the parallel nature of the dataset to project pre-annotations from the EN passages to the MSA passages as part of a bootstrapping approach. The goal was to maximize the alignment between the EN and the MSA spans that provide answers to the questions. Our semi-automated approach followed the following steps:
  1. 1.
    Marking as ‘X’ the MSA questions labelled ‘X’ in English to avoid re-annotation. In our experience, this approach proved consistent for both English and MSA.
  2. 2.
    Translation of the EN span to MSA using Helsinki MT tool Tiedemann and Thottingal (2020)
  3. 3.
    Using a sliding-window approach to identify the MSA span with the same length and highest n-gram overlap with the EN span.
  4. 4.
    Marking an approximate location of the EN span in the MSA passage
  5. 5.
    If no equivalent translation was found, an approximate MSA location based on the EN span’s word offset was used. [^2] An example of this case is when the span is 1-2 words long and its translation is not found in the MSA passage.
  6. 6.
    Native Arabic speakers reviewed the pre-annotated spans according to the MSA-based annotation guidelines and made corrections where necessary.
The 415 QA pairs were divided amongst four native Arabic speaking annotators, with the annotations subsequently reviewed by independent reviewers. The annotators were presented with the MSA Belebele passage, the original MSA Belebele question and a copy of the passage where the span was pre-annotated. During the review process, the annotators could either: leave the span unchanged or modify the span markers, and label as ‘BB’ where necessary, or label the QA as ‘X’ to be removed.
The MSA guidelines are generally similar to the EN guidelines except for containing Arabic examples instead of English. The MSA guidelines match EN in using white-space boundaries as the primary word delimiter. This unavoidably leads to a difference in span scope from EN since Arabic is a morphologically rich language with numerous clitics. As a result some proclitic prepositions, conjunctions, and the definite article, as well as pronominal enclitics are kept inside the spans.
4.6 Inter-Annotator Agreement
We carried out an inter-annotation (IAA) study in order to assess the level of agreement across annotators, and get an indication of how reliable the guidelines were. [^3] As the dataset was small and fully manually reviewed, the IAA study did not serve to evaluate the data quality. For the IAA study, we selected 50 English QA pairs for annotation by three separate annotators. The three annotations for each QA pair were then compared. Instead of calculating an F-score measure, which would give indication of quality of annotations, we utilize the $\gamma$ measure for Inter-Annotator Agreement and Alignment Mathet et al. (2015). [^4] Gamma ranges between 0 and 1 where a higher value indicates higher agreement between annotators. This measure involves identifying the optimal alignment among annotations provided by multiple annotators, aiming to align annotations with high similarity and compute the average dissimilarity between them. Specifically, the average dissimilarity is calculated by averaging dissimilarities among aligned annotations, considering both their categories and positions. The alignment with the minimum mean dissimilarity, selected from all potential alignments initially evaluated during computation, is employed. The final inter-annotator agreement ($\gamma$ measure) was computed on the three sets of 50 QA pairs yielding a high $\gamma$ value of $0.81$, indicating that the designed annotation guidelines are reliable. Given that the MSA annotation guidelines were based on the EN guidelines, we posit that their quality was just as high. The three sets of 50 EN IAA QAs were merged and reviewed to produce a final version for inclusion in our parallel dataset. See Appendix A for more details and visualization of the span combination process.
5 Experiments and Results
5.1 Experimental Setup
Datasets
We use two datasets in our experiments: the commonly used state-of-the art QA dataset, SQuAD Rajpurkar et al. (2016) and our Belebele-EQA, both of which are Wikipedia-based. For SQuAD, we only used the validation portion of the dataset which contains 10,657 English EQA pairs. For Belebele-EQA, we have two versions: (i) Belebele-EQA-All contains 329 QA pairs for EN, MSA and five dialects, and (ii) Belebele-EQA-Sub excludes 44 BB-annotated questions, leaving 285 QA pairs.
QA Evaluation Tool
To run the QA task on the above datasets, we used PrimeQA Sil et al. (2023), an open-source [^5] https://github.com/primeqa/primeqa tool that is built on top of the HuggingFace Wolf et al. (2020) library to support research on running different QA tasks. We used two existing models, namely PrimeQA/NQ+TyDi [^6] PrimeQA/nq_tydi-reader-xlmr_large-20221210 and PrimeQA/NQ+TyDi+SQuAD, [^7] PrimeQA/nq_tydi_sq1-reader-xlmr_large-20221110 which are 550m parameters XLM-RLarge Conneau et al. (2020) based models that were finetuned for the multilingual TyDi QA task Clark et al. (2020). Both models were finetuned on the Natural Questions dataset Kwiatkowski et al. (2019) and TyDi Clark et al. (2020). The latter, however, was further fine-tuned on the training portion of the SQuAD dataset.
Evaluation Metrics
We report the results using the commonly used F1-score and Exact Match (EM) score. F1-score is the harmonic mean of Precision and the Recall on word-level uni-grams. The default F1 and EM scores as implemented in PrimeQA toolkit preprocess references and predictions by dropping punctuation, extra white space, and for EN, lower casing and removing the articles a/an/the. We also introduce two normalized variants, F1n and EMn, where we normalize references and predictions by removing the respective language stopwords as defined in NLTK Bird and Loper (2004).
5.2 Results and Analysis
In this section, we present the evaluation results for the Belebele-EQA and SQuAD datasets.
SQuAD vs. BELEBELE-EQA
We evaluate the SQuAD dataset as well as the EN-EN QA pairs from the Belebele-EQA dataset using the PrimeQA toolkit (See Appendix B for details) This experiment allows us to (a) verify that the chosen pretrained models are capable of solving the QA task at-hand without further finetuning, and (b) establish a sense of difficulty for the Belebele-EQA datasets compared to the commonly used SQuAD. We observe the following based on the results in Table 2: First, on SQuAD, the F1 shows a negligible difference between the two pretrained models, while EM is surprisingly lower by 3.4% for the model that was finetuned on the SQuAD dataset. [^8] The state-of-the-art F1 and EM scores on SQuAD are 93.2 and 90.9, respectively, as per the public PapersWithCode leaderboard: https://paperswithcode.com/sota/question-answering-on-squad20 In contrast, the results on Belebele-EQA-All show that additional finetuning on SQuAD yielded better performance both in F1 and EM.
Second, the best Belebele-EQA (All and Sub) F1 scores (71.0 and 76.6) are 18.6 and 14 points less than SQuAD’s F1 score, respectively. Similarly, the best respective EM scores are 50.5 and 56.1, or 28.6 and 23 points less than that of the SQuAD dataset. Finally, and as expected, removing the Belebele-EQA hard questions (Belebele-EQA-Sub) resulted in better performance on both F1 and EM measures.
BELEBELE-EQA Results
Based on the baseline results presented above, we report in Table 3 on the different question language setups using the best identified model: PrimeQA/NQ+TyDi+SQuAD. For specific experiments, we refer to the pair of languages in passage-question, e.g., EN-EN, or MSA-Iraqi.
Monolingual Experiments (EN-EN, MSA-MSA)
The basic standard monolingual setups have the highest performing F-1 scores, with EN being higher than MSA by 9% absolute. Despite being a multilingual model, this performance difference is not unexpected given the relative resource richness of EN compared with MSA.
Dataset Questions F1 EM
PrimeQA/NQ+TyDi
SQuAD 10,570 90.5 82.5
Belebele-EQA-All 329 68.1 46.5
Belebele-EQA-Sub 285 71.3 50.5
PrimeQA/NQ+TyDi+SQuAD
SQuAD 10,570 90.6 79.1
Belebele-EQA-All 329 71.0 50.5
Belebele-EQA-Sub 285 76.6 56.1
Table 2: Evaluating the SQuAD and belebele-EQA datasets using the PrimeQA dataset. (EM: Exact Match. Scores are percentages. Higher is better.)
Passage Question All Sub
F1 EM F1n EMn F1 EM F1n EMn
EN EN 71.0 50.5 74.0 60.2 76.6 56.1 79.5 65.6
MSA 59.9 39.8 62.7 47.7 62.9 43.5 65.6 51.2
[…]
DA–Levantine 49.5 31.9 51.2 36.5 51.9 35.1 53.7 39.6
DA–Gulf 50.4 31.3 52.6 37.4 52.1 33.0 54.2 38.9
DA–Morrocan 36.3 18.5 38.5 21.0 37.0 20.0 39.3 22.1
DA–Egyptian 48.8 30.1 51.5 36.2 51.3 33.3 54.0 39.3
DA–Average 45.9 27.9 48.0 32.5 47.7 30.2 49.8 34.6
[…]
DA–Levantine (TEN) 65.4 46.2 68.2 53.5 69.7 50.2 72.3 57.9
DA–Gulf (TEN) 65.4 43.8 68.2 52.9 69.5 48.1 72.2 57.2
DA–Morrocan (TEN) 63.0 41.0 65.7 49.5 67.0 44.9 69.6 53.0
DA–Egyptian (TEN) 65.7 45.0 68.7 53.8 70.0 49.5 72.9 58.2
DA–Average (TEN) 63.9 43.0 66.7 51.3 68.1 47.1 70.8 55.4
MSA EN 60.4 41.6 62.3 46.8 64.5 46.0 66.7 50.9
MSA 62.0 40.7 64.3 46.2 65.6 44.2 68.1 49.5
[…]
DA–Levantine 54.2 35.3 55.7 38.6 58.5 39.3 60.4 43.2
DA–Gulf 56.6 36.2 58.4 40.7 59.3 39.6 61.3 44.2
DA–Morrocan 45.1 26.7 46.6 30.1 48.1 29.8 49.6 33.3
DA–Egyptian 54.5 34.0 56.6 39.2 57.3 37.5 59.5 42.5
DA–Average 52.1 32.9 53.7 36.7 55.5 36.5 57.3 40.5
[…]
DA–Levantine (TMSA) 58.5 38.3 60.3 43.2 62.1 42.1 64.1 46.7
DA–Gulf (TMSA) 57.9 38.0 59.9 42.9 61.0 40.7 63.3 46.0
DA–Morrocan (TMSA) 56.5 36.2 58.6 42.6 59.4 39.3 61.5 45.3
DA–Egyptian (TMSA) 58.5 37.7 60.4 42.2 62.1 41.1 64.2 46.0
DA–Average (TMSA) 56.6 36.3 58.5 41.2 59.7 39.4 61.8 44.4
Table 3: Monolingual and Cross-lingual evaluation of the Belebele-EQA dataset using the PrimeQA toolkit with PrimeQA/NQ+TyDi+SQuAD. (EM: Exact Match; F1n and EMn are the text-normalized versions of F1 and EM, respectively; DA: Dialectal Arabic; TEN/MSA: Question translated to EN/MSA; Scores are %s; Higher is better.)
Cross-Lingual Experiments (MSA-EN, EN-MSA)
The basic cross-lingual setups are next in rank of performance with MSA-EN F1 being only slightly better than EN-MSA (0.5% absolute). The performance drop due to question language is much higher for EN passages (11.1% absolute) compared to MSA passages (1.6% absolute). This suggests that the model struggles in cross-lingual settings in ways independent of its monolingual performance.
Dialectal Experiments (MSA-DA, EN-DA)
The DA questions consistently lead to lower performance compare to the above-mentioned settings, and with performance with MSA passages being better on average than on EN passages (by 6.3% absolute). The specific dialects seem to follow a common order of Gulf$>$(Egyptian|Levantine)$>$Iraqi$>$Moroccan, which we speculate may reflect the order of their question sets’ similarity to MSA. Automatically translating dialectal questions into the target passage language using Google Translate [^9] For DA-to-English, we translated using Google Translate (AR-to-EN), while for DA-to-MSA, we used English as an intermediate language (DA $\Rightarrow$ EN $\Rightarrow$ MSA): Google Translate (AR-to-EN $\Rightarrow$ EN-to-AR). results in notable improvements: an average increase of 18% for English passages and 4.4% for MSA passages (All setting). Moroccan saw the greatest benefit from this translation method, with F1 increases of 26.7% for EN and 11.4% for MSA.
BELEBELE-EQA All vs Sub
Across all experiments, the easier subset (Sub) has higher scores than (All): on average, 2.6% absolute F1 increase for EN passages, and 1.5% for MSA passages. The highest difference is for EN-EN (5.4% absolute F1) and the lowest for MSA-Gulf (0.6%).
Normalized F1 and EM
The normalization results are generally consistent in rank order; but the effect of normalization in EN is much higher than MSA: F1 (2.3% vs 1.7%) and EM (5.8% 4.3%). We note that the normalization process reduces the length of references and predictions by 30% for EN, but only 10% for MSA.
6 Discussion
Our motivation for this study was to create a practical and reliable multilingual EQA benchmark dataset, without the need to start from scratch or resort to creating synthetic data. Given our languages of interest, repurposing the Belebele MCQA dataset seemed like a feasible solution. However, answering the research question on how feasible this approach is involves considering the various insights and findings presented in this paper.
We found that while repurposing is possible, our exploratory study showed that it required significant manual effort. Contrary to initial assumptions, we discovered that the original answers from the multiple-choice question pairs were not suitable for automated approaches. In fact, the only multiple-choice types in Table 1 that resemble non-complex SQUAD-style QA pairs are (a) Exact.
Rather than offering a scalable solution here, instead our manual efforts and the subsequent insights serve to make this repurposing task much more feasible and potentially automated by other dataset developers. We can see that the repurposing process is much more feasible when the differences between the QA styles are fully understood. We found that, through an effective filtering process, we could easily eliminate many QA pairs that did not lend themselves easily to EQA. While this resulted in reducing the dataset size by more than half, this leaves open new avenues for examining how those ‘unsuitable’ QA pairs could be reformulated to be useful in EQA. Repurposing a multilingual dataset like Belebele means that its parallelism could be exploited through a bootstrapping approach. We have learned that once one language dataset is manually annotated and verified, the spans could be relatively easily projected to a new language from the wider dataset provided a translation system is available (see Section 4.5). Likewise, few changes were required when adapting our annotation guidelines from EN to MSA, suggesting that this adaptability may hold for other languages. Finally, our experimental results confirm that our approach produced a useful data set that can be used for benchmarking cross-lingual EQA for all Belebele languages.
7 Conclusion and Future Work
We repurposed the MCQA Belebele dataset to create a new EQA dataset for English and MSA, with cross-lingual benchmark results for dialectal Arabic QA. [^10] https://github.com/mbzuai-nlp/MultiChoice2ExtractiveQA We also provided guidelines for practitioners exploring QA in other Belebele languages. Our work, including negative insights, aims to enhance understanding of the challenges and opportunities in reformulating NLP tasks.
For future work, we plan to support repurposing Belebele-EQA for other languages, potentially automating the process. The dataset offers a resource for evaluating such methods. We also aim to use it for fine-tuning general-purpose Arabic QA models and improving EQA evaluation metrics, focusing on span length.
Limitations
We would like to point to some limitations of our study. First, while we are addressing a general problem, in our study, out of all the other Belebele under-resourced languages, we only focused on cross-lingual evaluation of five Arabic dialects. While we believe that our findings are more generally applicable, we would like to note that each language and dataset may present unique challenges and opportunities. Second, our repurposing was based on a single dataset, Belebele, and for a single task (extractive question answering in the style of machine reading comprehension), and we note that other tasks may require some specific considerations.
Ethics and Broader Impact
Data Bias.
It is important to acknowledge the possibility of biases within our dataset due to subjective human judgments, or due to the influence of the original task’s data (e.g. a narrow cultural representation across the questions).
Broader Impact and Potential Use.
Our repurposing insights presented here can help speed up the creation of resources and tools for resource-poor languages and dialects. We do not see any immediate threats from its use.
Appendix A Processing IAA Spans for Visualization
In Figure 1, we visualize the degree of agreement among the annotators on the 42 QA pairs that remained from the IAA study (See Section 4.6). We combine the annotation spans into one artificial document in order to visualize the results in a meaningful way where this process is explained later in this section.
In this figure, the x-axis represents the artificial document that is 450 words long. The vertical, colored lines represents a span in that document. The three different colors represent the three annotators and the number on the span is the question number. 42 out of the 50 QA pairs remained as eight questions were labelled as ‘X’ by all three annotators and were excluded. As annotators were allowed to label a questions as ‘X’, some questions have less than three annotations, e.g. Question 38.
Figure 1: Visualization of the IAA agreement for 42 questions. The colored lines indicate a span. Each color represents an annotator. The number on the colored line is a question ID.
A.1 Combining the Spans Into One Artificial Document.
While we could visualize the agreement between annotators for each QA pairs separately, this would have resulted in 42 different figures, one for each QA pair. Instead, we decided to combine the annotated spans from each annotator to create one large artificial document where the annotated spans would appear in different locations. One simple approach to this is combining the whole passages which will offset the location of the annotated span by the length of all the passages that came before it. This, however, will result in a huge document that is too long to have a meaningful visualization.
A.2 Compressing the Artificial Document.
To overcome this issue, we compressed the document by removing the large un-annotated text spans. Note that the annotation are technically a start and end positions in a document and our task is basically to measure (and visualize) the chosen start and end position of each QA pair across the three annotators. This means that the QA pairs are independent, and ordering them in the document or shifting their locations –as long as they do not overlap and they maintain their original length– should not affect them.
To illustrate, assume that annotator ($a_{1}$) for passage ($p_{1}$) chose the span (3, 7) indicating that the answer is from word ($w_{3}$) to ($w_{7}$) while annotator ($a_{2}$) chose the span (4, 10). We can see that ($a_{1}$) and ($a_{2}$) disagree in the regions: ($w_{4}$) and from ($w_{8}$) to ($w_{1}0$). If we add passage ($p_{0}$) which contains 20 words before ($p_{1}$) to make one document resulting in shifting all the words in ($p_{1}$), as well as in the spans ($a_{1}$) and ($a_{2}$) by 20 words, we see that the disagreement remains the same. After the shift, ($a_{1}$) and ($a_{2}$) change from (3, 7) and (4, 10) to (23, 27) and (24, 30), respectively. As can be noticed, the disagreement between ($a_{1}$) and ($a_{2}$) remained in two regions with the same length for each region, that is ($w_{2}4$) and ($w_{2}8$) to ($w_{3}0$).
Appendix B PrimeQA Experimental Details
The experiments discusses in Section 5.2 were conducted on an Apple M1 Max MacBook Pro with 32GBs of memory. Since we used the aforementioned models in inference mode, each experiment with 329 questions (E.g. the combination of En Passage and En Questions) took approximately 30 seconds to complete. Note that additional time may be required to download each model for the first time. The 56 core experiments (without translation) in Table 3 (2 Passages x 7 question types x 2 configurations x 2 normalization schemes) are expected to require 28 minutes of running.
Appendix C Arabic Examples of MCQ-Span Types
Table 4: This table shows examples in Arabic demonstrating part of the range of types of MCQs in Belebele, and the respective issues they pose for extractive QA in this work. These examples are parallel to the English examples in Table 1.
Appendix D Guidelines for English Answer Span Annotation
See pages 2-6 of BelebeleQA-Annotation-Guidelines_EN_Paper_Appendix.pdf
Appendix E Guidelines for Modern Standard Arabic Answer Span Annotation